OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 108, -18, 3, -1/2, ...

OpenStudy (tkhunny):

Too bad there are only four terms. If we actually had a clue about the WHOLE pattern, we could have a solution. 108/(-18) = ?? -18/3 = ?? 3/(-1/2) = ??